OpenStudy (anonymous):

@tomhue can you check my answer and give me some help? and i will give you a medal again :)

OpenStudy (anonymous):

Sure.

OpenStudy (anonymous):

i know n equals 6

OpenStudy (anonymous):

and i think 150 is supossed to go on the other side

OpenStudy (anonymous):

Myra earns 25 dollars per lawn. "Per" signals multiplication, so 25L where L is lawns. She already has 350. Add 25L to it.

OpenStudy (davidusa):

it is 6. yes 150 should be on other side.

OpenStudy (anonymous):

And she wants 500 dollars or more, meaning greater than or equal to or >=. 350 + 25L >= 500 25L >= 150

OpenStudy (davidusa):

500-350= 150. 150/25 per lawn = 6 lawns.

OpenStudy (davidusa):

ty tom for medal

OpenStudy (anonymous):

thanks both of you :) you could be teachers

OpenStudy (davidusa):

lol

OpenStudy (davidusa):

next question?

OpenStudy (anonymous):

oh right

OpenStudy (davidusa):

lol